# spacing=nonfrench

### v1

@@ Abstract
We provide a complete characterization of all finite-dimensional quantum states that saturate the Margolus–Levitin quantum speed limit at arbitrary Uhlmann–Jozsa fidelity. Employing a purification-based approach, we prove that mixed-state saturation occurs precisely when three structural criteria are fulfilled: the state’s support is confined to the sum of two energy eigenspaces (the ground level and a single excited level); each eigenvector of the state with nonzero weight is a fixed superposition of one ground- and one excited-state energy eigenvector (determined by the minimizer of the objective function identified by Giovannetti et al.) and all such eigenvectors evolve in mutually orthogonal two-dimensional subspaces. These requirements impose a strict rank bound, ruling out saturation by any faithful state. For qubit systems, we derive a purity-resolved and tight Margolus–Levitin bound that reduces to the pure-state result in the limit of unit purity. Through a time-reversal argument, we further extend the dual Margolus–Levitin quantum speed limit to mixed states and establish the corresponding saturation conditions.
@@ I  Introduction
This time–energy relation—commonly referred to as the Margolus–Levitin quantum speed limit—is a fundamental bound that complements the Mandelstam–Tamm time–energy uncertainty relation [2, 3, 4]. It plays a significant role in the conceptual foundations of quantum theory and in the quantitative analysis of dynamical processes central to emerging quantum technologies [5, 6, 7, 8, 9]. Applications span a wide range of areas, including performance limits in quantum computation [10, 11, 12], precision metrology [13, 15, 14], bounds on information-processing rates [16, 17, 18], the charging of quantum batteries [19, 20, 21], and quantum control and shortcuts to adiabaticity [22].
@@
The inequality in Eq. (2) can be derived directly for qubit systems, and, supported by numerical evidence, Giovannetti et al. conjectured that it should hold for arbitrary finite-dimensional systems as well. Remarkably, a complete analytical proof did not appear until two decades later [24]. This delay can be attributed in part to the fact that the Margolus–Levitin speed limit possesses a geometric character that differs fundamentally from that of other quantum speed limits, as discussed in Ref. [24].
@@
In this work, we identify all systems that saturate the Margolus–Levitin quantum speed limit, meaning those whose dynamics achieve equality in Eq. (2). While Ref. [24] provides a complete characterization of the pure states that attain the Margolus–Levitin bound, we extend this characterization to include mixed states as well (Secs. III and IV), thereby broadening the class of systems for which the bound is tight.
From this characterization, it follows that any state saturating the Margolus–Levitin quantum speed limit must have a highly constrained rank. In particular, its support is confined to the direct sum of two energy eigenspaces, and its dimension cannot exceed the smaller of the multiplicities of those eigenvalues. Consequently, faithful states can never saturate the Margolus–Levitin bound. This observation is made explicit in Sec. V, where we present a refined Margolus–Levitin quantum speed limit applicable to mixed—and therefore faithful—qubits.
In Sec. V, we also extend the dual Margolus–Levitin quantum speed limit derived by Ness et al. [27] to mixed states and characterize the systems that saturate this dual bound. This characterization follows directly from that obtained for systems saturating the standard Margolus–Levitin bound, since the dual speed limit can be recovered from the original via a time-reversal argument.
@@ II  The Margolus–Levitin speed limit
Consider a system in a state $\rho$ undergoing unitary evolution generated by a Hamiltonian $H$. Let $E=\operatorname{tr}(\rho H)$ denote the system’s expected energy, and let $E_{0}$ be the ground-state energy, defined as the smallest populated eigenvalue of $H$. The time $\tau$ required for the system to evolve into a state having Uhlmann–Jozsa fidelity $\delta$ with the initial state is then bounded from below by the Margolus–Levitin quantum speed limit:
@@
We say that an energy level, that is, an eigenvalue of $H$, is populated if the support of the state is not orthogonal to the corresponding eigenspace. Equivalently, the eigenvalue occurs with nonzero probability in an energy measurement. Because the Hamiltonian is time-independent, this probability is conserved. Consequently, the support of the state is contained in the sum of the eigenspaces associated with the populated energies.
@@
The Margolus–Levitin speed limit does not, in general, extend to systems governed by time-dependent Hamiltonians. As shown in Ref. [28], no canonical extension of the Margolus–Levitin bound exists in this setting.
@@ II.A  Purification and Uhlmann’s Theorem
Any other purification of $\rho$ has the form $\mathbbm{1}\otimes U|w\rangle$ for some unitary operator $U$ acting on $\mathcal{H}$. Uhlmann’s theorem [25] asserts that the fidelity between two mixed states equals the maximal possible overlap between their purifications. Specifically, if $|w_{1}\rangle$ and $|w_{2}\rangle$ are purifications of $\rho_{1}$ and $\rho_{2}$, then
@@ II.B  The Margolus–Levitin quantum speed limit for systems in mixed states
Figure 1: Graph of the function $\alpha$, which appears in the numerator of the Margolus–Levitin bound, shown as a function of the fidelity $\delta$ between the initial and final states. As evident from the plot, the function is strictly decreasing. A proof of this property is provided in Appendix A.
@@ III  Systems saturating the Margolus–Levitin quantum speed limit
In this section, we provide a complete characterization of all systems that saturate the Margolus–Levitin quantum speed limit. Reference [24] includes a full characterization of the systems in pure states that attain the Margolus–Levitin bound. We begin by recalling that result.
@@ III.A  Systems in pure states that saturate the Margolus–Levitin quantum speed limit
$|\psi\rangle=\sqrt{\frac{1-z_{\delta}}{2}}|E_{0}\rangle+\sqrt{\frac{1+z_{\delta}}{2}}|E_{1}\rangle,$ (12)
where $|E_{0}\rangle$ is a normalized ground-state vector, $|E_{1}\rangle$ is a normalized energy eigenvector corresponding to a higher eigenvalue $E_{1}$, and $z_{\delta}$ is the unique value of $z$ at which the objective function $f_{\delta}(z)$, defined in Eq. (4), attains its minimum. The existence and uniqueness of this value are established in Appendix B. Figure 2 plots $f_{\delta}(z)$ for several values of $\delta$, with the associated minimizers indicated along the horizontal axis.
@@
This characterization follows from the observation that the system must effectively behave as a qubit, that is, evolve within the span of two eigenvectors of the Hamiltonian. This was shown in Ref. [24]. For pure qubit systems, it is then straightforward to determine which states saturate the Margolus–Levitin bound. An explicit proof that $|\psi\rangle$ can be written in the form of Eq. (12) is provided in Sec. V, which analyzes qubit systems in detail.
@@ III.B  Systems in mixed states that saturate the Margolus–Levitin quantum speed limit
    The support of $\rho$ is contained within the direct sum of two eigenspaces of $H$.
Since the populations of the energy levels are conserved, the system effectively behaves as a two-level system, possibly with degenerate levels. Let the corresponding energies be $E_{0}$ and $E_{1}$, with $E_{0}$ denoting the lower of the two.
@@
    Every pair of orthogonal eigenvectors of $\rho$ with nonzero eigenvalues evolves within orthogonal two-dimensional subspaces of the Hilbert space.
This property implies that the rank of the state cannot exceed the smaller of the multiplicities of $E_{0}$ and $E_{1}$. Consequently, a faithful state can never saturate the Margolus–Levitin bound. In Sec. V, we refine the Margolus–Levitin speed limit for qubit systems in a manner that remains tight irrespective of the state’s purity.
@@
$|\psi_{j}\rangle=\sqrt{\frac{1-z_{\delta}}{2}}|E_{0}^{j}\rangle+\sqrt{\frac{1+z_{\delta}}{2}}|E_{1}^{j}\rangle,$ (14)
where $|E_{0}^{j}\rangle$ and $|E_{1}^{j}\rangle$ are normalized eigenvectors of $H$ associated with the eigenvalues $E_{0}$ and $E_{1}$, respectively. These energy eigenvectors are pairwise orthogonal,
$\langle E_{0}^{j}|E_{0}^{k}\rangle=\langle E_{1}^{j}|E_{1}^{k}\rangle=\delta_{jk},\quad\langle E_{0}^{j}|E_{1}^{k}\rangle=0,$ (15)
@@
$E=\bigg(\frac{1-z_{\delta}}{2}\bigg)E_{0}+\bigg(\frac{1+z_{\delta}}{2}\bigg)E_{1},$ (16)
@@ IV.C  Properties (i)–(iii) are necessary
The first equality in Eq. (11) then implies that the Margolus–Levitin bound for pure states is saturated by $|w\rangle\langle w|$. From the characterization in Sec. III, $|w\rangle$ must therefore lie in the direct sum of the eigenspaces of $H\otimes\mathbbm{1}$ associated with $E_{0}$ and some higher eigenvalue $E_{1}$. These eigenspaces are $\mathcal{E}_{0}\otimes\mathcal{H}$ and $\mathcal{E}_{1}\otimes\mathcal{H}$, where $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$ are the eigenspaces of $H$ corresponding to $E_{0}$ and $E_{1}$. Since
$|w\rangle\in(\mathcal{E}_{0}\otimes\mathcal{H})\oplus(\mathcal{E}_{1}\otimes\mathcal{H})=(\mathcal{E}_{0}\oplus\mathcal{E}_{1})\otimes\mathcal{H},$ (19)
the support of $\rho$ is contained in the sum of $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$:
$\operatorname{supp}\rho=\operatorname{supp}\operatorname{tr}_{2}|w\rangle\langle w|\subset\mathcal{E}_{0}\oplus\mathcal{E}_{1}.$ (20)
@@
In Appendix D, we show that condition (22) implies that the compression of $U_{\tau}$ to the support of $\rho$ is proportional to the orthogonal projection onto that support, and that the proportionality constant has modulus $\sqrt{\delta}$. Thus, if $P$ denotes the orthogonal projection onto the support of $\rho$, condition (22) yields
@@
Let $|\psi_{1}\rangle=|\psi\rangle$ and $p_{1}=p$, and expand $\rho$ in a spectral decomposition as in Eq. (13), where $r$ is the rank of $\rho$, so that all $p_{j}>0$ and each $|\psi_{j}\rangle$ lies in the support of $\rho$. We now show that all states $|\psi_{j}\rangle\langle\psi_{j}|$ have the same expected energy. Since each of these states is supported in $\mathcal{E}_{0}\oplus\mathcal{E}_{1}$ and is nonstationary, it must populate the ground-energy eigenspace. Applying the Margolus–Levitin bound together with Eq. (24) yields, for every $|\psi_{j}\rangle$,
@@
Let $P_{0}$ and $P_{1}$ be the orthogonal projections onto the eigenspaces $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$, respectively. Since the support of $\rho$ is contained in $\mathcal{E}_{0}\oplus\mathcal{E}_{1}$, Eq. (23) implies that
$e^{-i\tau E_{0}}PP_{0}P+e^{-i\tau E_{1}}PP_{1}P=\sqrt{\delta}\,e^{i\theta}P.$ (28)
Together with $PP_{0}P+PP_{1}P=P$, this shows that the compressions of the eigenspace projectors to the support of $\rho$ are each proportional to $P$:
$PP_{0}P$ $=q_{0}P,\quad$ $q_{0}=\frac{\sqrt{\delta}\,e^{i\theta}-e^{-i\tau E_{1}}}{e^{-i\tau E_{0}}-e^{-i\tau E_{1}}},$ (29a)
$PP_{1}P$ $=q_{1}P,\quad$ $q_{1}=\frac{\sqrt{\delta}\,e^{i\theta}-e^{-i\tau E_{0}}}{e^{-i\tau E_{1}}-e^{-i\tau E_{0}}}.$ (29b)
@@
$q_{1}\delta_{jk}$ $=\langle\psi_{k}|P_{1}|\psi_{j}\rangle=\frac{1}{2}(1+z_{\delta})\langle E_{1}^{j}|E_{1}^{k}\rangle.$ (30b)
These identities show that the conditions in Eq. (15) are satisfied and hence that the eigenvectors $|\psi_{j}\rangle$ evolve within mutually orthogonal two-dimensional subspaces.

### v2

@@ Abstract
We provide a complete characterization of all finite-dimensional quantum systems that saturate the Margolus–Levitin quantum speed limit at arbitrary Uhlmann–Jozsa fidelity. Employing a purification-based approach, we prove that mixed-state saturation occurs precisely when three structural criteria are fulfilled: the state’s support is confined to the sum of two energy eigenspaces (the ground level and a single excited level); each eigenvector of the state with nonzero weight is a fixed superposition of one ground- and one excited-state energy eigenvector (determined by the minimizer of the objective function identified by Giovannetti et al.) and all such eigenvectors evolve in mutually orthogonal subspaces. These requirements impose a strict rank bound, ruling out saturation by any faithful state. For quantum bits, we derive a purity-resolved and tight Margolus–Levitin bound that reduces to the pure-state result in the limit of unit purity. Through a time-reversal argument, we further extend the dual Margolus–Levitin quantum speed limit to mixed states and establish the corresponding saturation conditions.
@@ I  Introduction
This time–energy relation, commonly referred to as the Margolus–Levitin quantum speed limit, is a fundamental bound that complements the Mandelstam–Tamm time–energy uncertainty relation [2, 3, 4]. It plays a significant role in the conceptual foundations of quantum theory and in the quantitative analysis of dynamical processes central to emerging quantum technologies [5, 6, 7, 8, 9]. Applications span a wide range of areas, including performance limits in quantum computation [10, 11, 12], precision metrology [13, 15, 14], bounds on information-processing rates [16, 17, 18], the charging of quantum batteries [19, 20, 21], and quantum control and shortcuts to adiabaticity [22].
@@
The inequality in Eq. (2) can be derived directly for qubits, and, supported by numerical evidence, Giovannetti et al. conjectured that it should hold for arbitrary finite-dimensional systems as well. Remarkably, a complete analytical proof did not appear until two decades later [24]. This delay can be attributed in part to the fact that the Margolus–Levitin speed limit possesses a geometric character that differs fundamentally from that of other quantum speed limits, as discussed in Ref. [24].
@@
In this work, we identify all systems that saturate the Margolus–Levitin quantum speed limit, that is, those whose dynamics achieve equality in Eq. (2). Specifically, for any finite-dimensional system evolving under a time-independent Hamiltonian, we derive necessary and sufficient conditions on the initial state ensuring that it evolves into a state of prescribed fidelity in a time saturating the Margolus–Levitin bound. While Ref. [24] provides a complete characterization of the pure states attaining the Margolus–Levitin bound, we extend this characterization to mixed states (Secs. III and IV), thereby broadening the class of states for which the bound is tight.
From this characterization, it follows that any state saturating the Margolus–Levitin quantum speed limit must have a highly constrained rank. In particular, its support is confined to the sum of two energy eigenspaces, and its dimension cannot exceed the smaller of the multiplicities of those eigenvalues. Consequently, faithful states can never saturate the Margolus–Levitin bound. This observation is made explicit in Sec. V, where we present a refined Margolus–Levitin quantum speed limit applicable to mixed—and therefore faithful—quantum bits.
In Sec. V, we also extend the dual Margolus–Levitin quantum speed limit derived by Ness et al. [27] to mixed states and characterize the states for which this bound is saturated. This characterization follows directly from that obtained for systems saturating the standard Margolus–Levitin bound, since the dual speed limit can be recovered from the original via a time-reversal argument.
@@ II  The Margolus–Levitin speed limit
Consider a system in a state $\rho$ undergoing unitary evolution generated by a Hamiltonian $H$. Let $E=\operatorname{tr}(\rho H)$ denote the system’s expected energy, and let $E_{0}$ be the ground-state energy, defined as the smallest populated eigenvalue of $H$ (see Remark 1 below). The time $\tau$ required for the system to evolve into a state having Uhlmann–Jozsa fidelity $\delta$ with the initial state is then bounded from below by the Margolus–Levitin quantum speed limit:
@@
An energy level (an eigenvalue of $H$) is populated if the support of the state is not orthogonal to the corresponding eigenspace; equivalently, the eigenvalue occurs with nonzero probability in an energy measurement. For a time-independent Hamiltonian, these probabilities are conserved, so the state’s support remains within the sum of the populated eigenspaces.
In the literature on the Margolus–Levitin quantum speed limit, $E_{0}$ is typically defined as the smallest eigenvalue of $H$. If this eigenvalue is unpopulated, the smallest populated eigenvalue can be used instead, yielding a strictly larger lower bound on the evolution time. In this work, we take $E_{0}$ to be the smallest populated eigenvalue.
@@
The Margolus–Levitin speed limit does not, in general, extend to systems with time-dependent Hamiltonians. As shown in Ref. [28], no canonical extension of the Margolus–Levitin bound exists in this setting.
@@ II.A  Purification and Uhlmann’s theorem
Any other purification of $\rho$ has the form $\mathbbm{1}\otimes U|w\rangle$ for some unitary operator $U$ acting on $\mathcal{H}$. Uhlmann’s theorem [25] asserts that the fidelity between two mixed states equals the maximal possible overlap between their purifications. Specifically, if $|w_{1}\rangle$ and $|w_{2}\rangle$ are purifications of $\rho_{1}$ and $\rho_{2}$, respectively, then
@@ II.B  The Margolus–Levitin quantum speed limit for systems in mixed states
Figure 1: Plot of the function $\alpha$, which appears in the numerator of the Margolus–Levitin bound, shown as a function of the fidelity $\delta$ between the initial and final states. As evident from the plot, the function is strictly decreasing. A proof of this property is provided in Appendix A.
@@ III  Systems saturating the Margolus–Levitin quantum speed limit
In this section, we derive necessary and sufficient conditions on a system’s state guaranteeing saturation of the Margolus–Levitin quantum speed limit. Reference [24] provides a complete characterization of the pure states for which the system attains the bound. We begin by recalling that result.
@@ III.A  Systems in pure states that saturate the Margolus–Levitin quantum speed limit
$|\psi\rangle=\sqrt{\frac{1-z_{\delta}}{2}}|E_{0}\rangle+\sqrt{\frac{1+z_{\delta}}{2}}|E_{m}\rangle,$ (12)
where $|E_{0}\rangle$ is a normalized ground-state vector, $|E_{m}\rangle$ is a normalized energy eigenvector corresponding to a higher eigenvalue $E_{m}$, and $z_{\delta}$ is the unique value of $z$ at which the objective function $f_{\delta}(z)$, defined in Eq. (4), attains its minimum. The existence and uniqueness of this value are established in Appendix B. Figure 2 plots $f_{\delta}(z)$ for several values of $\delta$, with the associated minimizers indicated along the horizontal axis.
@@
This characterization follows from the observation that the system must effectively behave as a qubit, that is, evolve within the span of two eigenvectors of the Hamiltonian. This was shown in Ref. [24]. For pure qubits, it is then straightforward to determine which states saturate the Margolus–Levitin bound. An explicit proof that $|\psi\rangle$ can be written in the form of Eq. (12) is provided in Sec. V, which analyzes qubits in detail.
@@ III.B  Systems in mixed states that saturate the Margolus–Levitin quantum speed limit
    The support of $\rho$ is contained within the sum of two eigenspaces of $H$.
Since the populations of the energy levels are conserved, the system effectively reduces to a two-level system, possibly with degenerate levels. The lower of the two populated energies is $E_{0}$, and we denote the other energy by $E_{m}$.
@@
    Any two orthogonal eigenvectors of $\rho$ with nonzero eigenvalues evolve within orthogonal two-dimensional subspaces of the Hilbert space.
This property implies that the rank of the state cannot exceed the smaller of the multiplicities of $E_{0}$ and $E_{m}$. Consequently, a faithful state can never saturate the Margolus–Levitin bound. In Sec. V, we refine the Margolus–Levitin speed limit for qubits in a manner that remains tight irrespective of the state’s purity.
@@
$|\psi_{j}\rangle=\sqrt{\frac{1-z_{\delta}}{2}}|E_{0}^{j}\rangle+\sqrt{\frac{1+z_{\delta}}{2}}|E_{m}^{j}\rangle,$ (14)
where $|E_{0}^{j}\rangle$ and $|E_{m}^{j}\rangle$ are normalized eigenvectors of $H$ associated with the eigenvalues $E_{0}$ and $E_{m}$, respectively. These energy eigenvectors are pairwise orthogonal,
$\langle E_{0}^{j}|E_{0}^{k}\rangle=\langle E_{m}^{j}|E_{m}^{k}\rangle=\delta_{jk},\quad\langle E_{0}^{j}|E_{m}^{k}\rangle=0,$ (15)
@@
$E=\bigg(\frac{1-z_{\delta}}{2}\bigg)E_{0}+\bigg(\frac{1+z_{\delta}}{2}\bigg)E_{m},$ (16)
@@ IV.C  Properties (i)–(iii) are necessary
The first equality in Eq. (11) then implies that the Margolus–Levitin bound for pure states is saturated by $|w\rangle\langle w|$. From the characterization in Sec. III, $|w\rangle$ must therefore lie in the sum of the eigenspaces of $H\otimes\mathbbm{1}$ associated with $E_{0}$ and some higher eigenvalue $E_{m}$. These eigenspaces are $\mathcal{E}_{0}\otimes\mathcal{H}$ and $\mathcal{E}_{m}\otimes\mathcal{H}$, where $\mathcal{E}_{0}$ and $\mathcal{E}_{m}$ are the eigenspaces of $H$ corresponding to $E_{0}$ and $E_{m}$. Since
$|w\rangle\in(\mathcal{E}_{0}\otimes\mathcal{H})\oplus(\mathcal{E}_{m}\otimes\mathcal{H})=(\mathcal{E}_{0}\oplus\mathcal{E}_{m})\otimes\mathcal{H},$ (19)
the support of $\rho$ is contained in the sum of $\mathcal{E}_{0}$ and $\mathcal{E}_{m}$:
$\operatorname{supp}\rho=\operatorname{supp}\operatorname{tr}_{2}|w\rangle\langle w|\subset\mathcal{E}_{0}\oplus\mathcal{E}_{m}.$ (20)
@@
In Appendix D, we show that the condition in Eq. (22) implies that the compression of $U_{\tau}$ to the support of $\rho$ is proportional to the orthogonal projection onto that support, and that the proportionality constant has modulus $\sqrt{\delta}$. Thus, if $P$ denotes the orthogonal projection onto the support of $\rho$, Eq. (22) yields
@@
Let $|\psi_{1}\rangle=|\psi\rangle$ and $p_{1}=p$, and expand $\rho$ in a spectral decomposition as in Eq. (13), where $r$ is the rank of $\rho$, so that all $p_{j}>0$ and each $|\psi_{j}\rangle$ lies in the support of $\rho$. We now show that all states $|\psi_{j}\rangle\langle\psi_{j}|$ have the same expected energy. Since each of these states is supported in $\mathcal{E}_{0}\oplus\mathcal{E}_{m}$ and is nonstationary, it necessarily populates the ground-energy level. Applying the Margolus–Levitin bound together with Eq. (24) yields, for every $|\psi_{j}\rangle$,
@@
Let $P_{0}$ and $P_{m}$ be the orthogonal projections onto the eigenspaces $\mathcal{E}_{0}$ and $\mathcal{E}_{m}$, respectively. Since the support of $\rho$ is contained in $\mathcal{E}_{0}\oplus\mathcal{E}_{m}$, Eq. (23) implies that
$e^{-i\tau E_{0}}PP_{0}P+e^{-i\tau E_{m}}PP_{m}P=\sqrt{\delta}\,e^{i\theta}P.$ (28)
Together with $PP_{0}P+PP_{m}P=P$, this shows that the compressions of the eigenspace projectors to the support of $\rho$ are each proportional to $P$:
$PP_{0}P$ $=q_{0}P,\quad$ $q_{0}=\frac{\sqrt{\delta}\,e^{i\theta}-e^{-i\tau E_{m}}}{e^{-i\tau E_{0}}-e^{-i\tau E_{m}}},$ (29a)
$PP_{m}P$ $=q_{m}P,\quad$ $q_{m}=\frac{\sqrt{\delta}\,e^{i\theta}-e^{-i\tau E_{0}}}{e^{-i\tau E_{m}}-e^{-i\tau E_{0}}}.$ (29b)
@@
$q_{m}\delta_{jk}$ $=\langle\psi_{k}|P_{m}|\psi_{j}\rangle=\frac{1}{2}(1+z_{\delta})\langle E_{m}^{j}|E_{m}^{k}\rangle.$ (30b)
Since $z_{\delta}\neq\pm 1$, these identities show that the conditions in Eq. (15) are satisfied and hence that the eigenvectors $|\psi_{j}\rangle$ evolve within mutually orthogonal two-dimensional subspaces.